19 Apr 18
The crew awoke in superb spirits. We conducted additional map training and MTU coordinate system training as well as topographical map indoctrination.
EVA team one left at 0800 with a compass and copy of map to ensure walking directions. Remaning crew did a full field day of sciemce HAB and it is standing by for inspection and for Korean film crew to use as staging area…as directed.
Additionally the stay behind team field day-ed lower portion of HAB and bathroom area. Finally we transferred ~450 gallons from 2nd storage tank to primary (black tank) leaving ~100 gallons in 2nd tank and black tank full. Also cleaned the water pump and pump container.
Inventoried, organized and cleaned 30 headphone / earpieces. Outatanding are 5 currently in use by our crew.
Afternoon EVA worked hard to remove water tank from the RAM and secure to MARS society trailer as mission complete on cleaning. They also cleaned RAM. The doctor recommends a test kit of sorts to ensure cleanliness prior to use and water consumption from container.
EVA CREW was a little tired / under the weather so they retreated to the HAB to do some training.
Did an inventory of all food and began prepping final report for end of mission.
IMPRESSED: Work ethic and “can do” spirit award goes to Andreea (Blue 03) who worked tirelessly on clean up, garden, bathroom and water transfer. BZ Blue 03.
POA&M: Intend on going out on EVAs tomorrow for final data collection and coming out of SIM not later than friday evening in preparation for cleaning and turnover wirh Gold crew.
Joseph Dituri, PhD(c)
CDR, U.S. Navy Diving Officer (ret)
Director – IBUM